# FedCQA: Answering Complex Queries on Multi-Source Knowledge Graphs via Federated Learning

### v2

@@ Abstract.
Complex logical query answering is a challenging task in knowledge graphs (KGs) that has been widely studied. The ability to perform complex logical reasoning is essential and supports various graph reasoning-based downstream tasks, such as search engines. Recent approaches are proposed to represent KG entities and logical queries into embedding vectors and find answers to logical queries from the KGs. However, existing proposed methods mainly focus on querying a single KG and cannot be applied to multiple graphs. In addition, directly sharing KGs with sensitive information may incur privacy risks, making it impractical to share and construct an aggregated KG for reasoning to retrieve query answers. Thus, it remains unknown how to answer queries on multi-source KGs. An entity can be involved in various knowledge graphs and reasoning on multiple KGs and answering complex queries on multi-source KGs is important in discovering knowledge cross graphs. Fortunately, federated learning is utilized in knowledge graphs to collaboratively learn representations with privacy preserved. Federated knowledge graph embeddings enrich the relations in knowledge graphs to improve the representation quality. However, these methods only focus on one-hop relations and cannot perform complex reasoning tasks. In this paper, we apply federated learning to complex query-answering tasks to reason over multi-source knowledge graphs while preserving privacy. We propose a Federated Complex Query Answering framework (FedCQA), to reason over multi-source KGs avoiding sensitive raw data transmission to protect privacy. We conduct extensive experiments on three real-world datasets and evaluate retrieval performance on various types of complex queries. The experimental results show that our proposed methods can effectively retrieve answers to cross knowledge graph queries while keeping raw data secret in local devices.
privacy preserving, knowledge graph, complex query answering, federated Learning
@@ 1. Introduction
Knowledge graphs (KG) are structured knowledge databases representing external knowledge using triplets and describing the relationships between entities. KG plays an important role and has been widely applied in numerous real-world applications, such as recommender systems (Wang et al., 2019; Cao et al., 2019) and search engines (Xiong et al., 2017). Among them, complex query answering (CQA) is an important but challenging task in knowledge graph reasoning and can be used for supporting various downstream tasks (Bai et al., 2023c; Ren et al., 2023). CQA aims to retrieve answers that satisfy given logical expressions (Hamilton et al., 2018; Ren et al., 2020), which are often defined in predicate logic forms with relation projection operations, existential quantifiers $\exists$, logical conjunctions $\wedge$, disjunctions $\vee$, etc. For example in Figure 1, given a logical query $q$, We aim to find all the research topic entities $V_{?}$ for which there exists Nobel Prize winners $V$ who were born in Germany and conducted studies in that specific field. CQA is challenging due to the incompleteness of knowledge graphs and the complexity of query matching (Bai et al., 2023c). To solve these problems, query encoding methods are commonly used in CQA. These methods try to parse complex queries into computation graphs and parameterize both queries and entities using different neural networks simultaneously. After the encoding process, these methods retrieve answers according to the similarity scores between the parameterized embeddings of answer sets and queries  (Hamilton et al., 2018; Ren et al., 2020).
@@
Existing query encoding methods have gained remarkable success in addressing complex query answering tasks. However, these query answering models are limited to the utilization of a single central knowledge graph and cannot be extended to multiple KGs. With the development of knowledge graphs, KGs have grown significantly in size and encompass knowledge from various domains. Consequently, constructing a knowledge graph that includes all related entities and relations has become difficult and it is impractical to access a central KG with all the triples needed (Peng et al., 2021a; Chen et al., 2021; Zhang et al., 2022). Collaborations between various knowledge graph holders are helpful to answer more complicated queries. For example, as shown in Figure 1, there are multiple knowledge graphs from different domains. A complex query may consist of entities and relations from multiple knowledge graphs, preventing a local query answering model on a single knowledge graph from answering that cross-graph queries. However, there are various reasons hindering the data sharing between knowledge graph holders. For example, the growing attention on privacy, regulations such as the General Data Protectioan Regulation (GDPR), and commercial interest between data holders, etc.
@@
Figure 1. An example of cross knowledge graph queries on distributed knowledge graphs. The relations and entities in a KG complex query can be from multiple knowledge graphs which cannot be solved in a single local knowledge graph.
@@
To solve the above challenges, federated learning (McMahan et al., 2017; Yang et al., 2019) has been proposed which allows multiple participants to distributedly train a global model collaboratively without raw data transmission. Federated learning has been widely applied in various domains (Long et al., 2020), including knowledge graph embedding. In such systems, raw knowledge graph triplets are kept on the local devices, participants train local models, and only gradient will be transferred to learn a global knowledge graph embedding model (Chen et al., 2021; Zhang et al., 2022; Peng et al., 2021a). In the training process, homomorphic encryption (HE) (Paillier, 1999), secure multi-party computation (SMPC) (Mohassel and Zhang, 2017), and differential privacy (Dwork, 2008; Geyer et al., 2017) are widely applied in federated learning to improve security. FedAvg (McMahan et al., 2017) is a commonly used technique in federated learning that updates global models by averaging local models trained on individual data and has been widely studied in various areas, including federated knowledge graph embedding (Chen et al., 2021; Zhang et al., 2022). Although federated learning can protect raw data, recent studies have indicated that learned representations can still potentially leak privacy (Hu and Song, 2023; Duddu et al., 2020), where an attacker can infer sensitive information from the learned embeddings.
@@
While federated learning has been widely applied in knowledge graph embeddings, existing federated knowledge graph embedding works only focus on learning high quality representations for simple downstream tasks, such as knowledge graph completion (Chen et al., 2021; Zhang et al., 2022; Huang et al., 2022; Tang et al., 2023), and lacks the ability of reasoning over knowledge graphs and retrieving answers to complex queries. To solve the constraints, we propose a Federated Complex Query Answering framework (FedCQA), to reason over multi-source KGs avoiding sensitive raw data transmission to protect data privacy. FedCQA can be applied to different central complex query answering models. It leverages federated learning techniques to train local query answering models in local knowledge graphs and align knowledge graph embeddings for global queries. Different from other federated knowledge graph embedding models, FedCQA server not only takes the responsibility of aggregating global models, but decomposing given complex queries to sub-queries to compute the query encoding and retrieve answers from distributed knowledge graphs under the protection of multi-party computation to avoid global model storage. Meanwhile, to better evaluate federated complex query answering systems’ performance, we create a benchmark on three widely used knowledge graph datasets. We evaluate our proposed FedCQA on the benchmark and the experiment results show the effectiveness of retrieving answers to complex queries from multi-source knowledge graphs. We summarize our major contributions as follows:
@@
  • To the best of our knowledge, we are the first to extend federated knowledge graph embedding systems to complex query answering tasks, which is critical for knowledge graph holders’ collaboration.
@@
  • Based on three public datasets, we propose a benchmark for evaluating the retrieval performance of federated complex query answering systems. The benchmark systematically evaluates the retrieval performance facing cross knowledge graph queries.
@@
  • We propose FedCQA, a federated complex query answering system that can retrieve answers to complex queries from distributed knowledge graphs with privacy preserved. Extensive experiments conducted on three datasets demonstrate its high performance facing cross knowledge graphs queries.
The rest of the paper is organized as follows. We review the related works in Section 2. Section 3 introduces the preliminary and definition of the distributed complex query answering systems. Section 4 introduces the framework of FedCQA in detail. Section 5 evaluates the performance of FedCQA on the benchmark based on three real-world datasets. Finally, we make a conclusion of our work in Section 6.
@@ 2. Related Work
In this section, we summarize the related work. Our work is closely related to Complex query Answering (CQA) and federated knowledge graph embedding.
2.1. Complex Query Answering
@@
Complex query answering (CQA) is a crucial task in knowledge graphs, where a model is learned to process and answer complex logical queries based on knowledge graph reasoning. Query encoding is effective and efficient for complex query answering. These methods represent complex queries into various structures and effectively search answers among candidate knowledge graph entities. GQE (Hamilton et al., 2018), Q2B (Ren et al., 2020) and HypeE (Liu et al., 2021) encode queries to vectors, hyper-rectangle and hyperbolic embeddings, respectively. To support negation operators, various encoding methods are proposed: Q2P (Bai et al., 2022) and ConE (Zhang et al., 2021) use multiple vectors to represent queries. BetaE (Ren and Leskovec, 2020), GammaE (Yang et al., 2022), PERM (Choudhary et al., 2021) propose to use various probabilistic distributions to encode complex logic graph queries. Some methods, like LMPNN (Wang et al., 2022) CQD (Arakelyan et al., 2020), Var2Vec (Wang et al., 2023a) take pre-train embeddings on simple link prediction tasks and apply logic operators to answer complex queries. Meanwhile, Neural structures are utilized to encode complex queries: BiQE (Kotnis et al., 2021) and KgTransformer (Liu et al., 2022) are proposed to use transformer, SQE (Bai et al., 2023c) applies sequential encoders, GNN-QE (Zhu et al., 2022) and StarQE (Alivanistos et al., 2021) use message-passing graph neural networks to encode queries, respectively. There are also encoding methods proposed to encode various knowledge graph types: NRN (Bai et al., 2023b) is proposed to encode numerical values, MEQE (Bai et al., 2023a) extends logical queries over events, states, and activities.
@@
While there are numerous existing complex query answering methods, these methods mainly focus on a single large knowledge graph. There are some methods proposed to reason over multi-view and temporal and varying knowledge graphs: MORA (Xi et al., 2022) ensembles multi-view knowledge graphs to scale up complex query answering, TTransE (Leblay and Chekol, 2018) and TRESCAL (Wang et al., 2015) can be applied on temporal knowledge graphs. However, these methods need raw data transmission and privacy protection is not considered. Conducting privacy-preserving complex query answering on multi-source knowledge graphs is still unexplored. With the growing attention on privacy and data protection, sensitive data is unable to freely circulate among data holders, and complex query answering is forced to be conducted on multiple knowledge graphs collaboratively. Our research introduces federated learning to existing complex query answering so that we can apply reasoning over distributed knowledge graphs without raw data sharing.
2.2. Federated Knowledge Graph Embedding
@@
In recent years, federated learning has emerged as a promising approach to address privacy and scalability concerns in machine learning. It allows data owners to participate in model co-construction without raw data transmission to reduce privacy leakage risks (McMahan et al., 2017; Yang et al., 2019). Various studies have been conducted to explore the potential of federated learning in different domains, such as recommender system (Yang et al., 2020), finance (Long et al., 2020) etc. Federated knowledge graph embedding is one of the domains in which federated learning has been widely applied. FedE (Chen et al., 2021) learns knowledge graph embeddings locally and aggregates all local models in a global server for higher representation quality. FedR (Zhang et al., 2022) proposes to learn representation with privacy-preserving relation aggregation to avoid privacy leakage risks in entity embedding and reduce communication costs. FedCKE (Huang et al., 2022), FedMKGC (Tang et al., 2023) extend federated learning to learn global representations from different domains and multilingual knowledge graphs. FedEC (Chen et al., 2022b) applies contrastive learning to tackle data heterogeneity in knowledge graphs. MaKEr (Chen et al., 2022a) and MorseE (Chen et al., 2022c) utilize meta-learning to transfer knowledge among knowledge graphs to train graph neural networks for unseen knowledge extrapolation and inductive learning. DP-FLames (Hu et al., 2023) quantifies the privacy threats and incorporates private selection in federated knowledge embeddings. FLEST (Wang et al., 2023b) decomposes the embedding matrix and enables the sharing of latent representations to reduce the risks of privacy leakage and communication costs, FedM (Hu et al., 2022b) splits the duty of aggregating entities and relations to reduce the risks of graph reconstruction attacks. FKGE (Peng et al., 2021b) applies differential privacy and avoids the need for a central server.
While existing federated knowledge graph embedding methods are proposed to distributedly learn high quality representations with privacy preservation, there are some works indicate that the learned embeddings are informative and vulnerable to various privacy attacks (Hu et al., 2022a), even in federated scenarios (Hu and Song, 2023; Hu et al., 2023). Besides, these methods all lack the ability to answer complex queries on multi-source knowledge graphs which is critical for more complicated downstream tasks. Our research expands the simple federated knowledge graph embedding to answer complex queries on distributed knowledge graphs.

### v3

# Federated Neural Graph Databases

@@ Abstract.
The increasing demand for large-scale language models (LLMs) has highlighted the importance of efficient data retrieval mechanisms. Neural graph databases (NGDBs) have emerged as a promising approach to storing and querying graph-structured data in neural space, enabling the retrieval of relevant information for LLMs. However, existing NGDBs are typically designed to operate on a single graph, limiting their ability to reason across multiple graphs. Furthermore, the lack of support for multi-source graph data in existing NGDBs hinders their ability to capture the complexity and diversity of real-world data. In many applications, data is distributed across multiple sources, and the ability to reason across these sources is crucial for making informed decisions. This limitation is particularly problematic when dealing with sensitive graph data, as directly sharing and aggregating such data poses significant privacy risks. As a result, many applications that rely on NGDBs are forced to choose between compromising data privacy or sacrificing the ability to reason across multiple graphs. To address these limitations, we propose Federated Neural Graph Database (FedNGDB), a novel framework that enables reasoning over multi-source graph-based data while preserving privacy. FedNGDB leverages federated learning to collaboratively learn graph representations across multiple sources, enriching relationships between entities and improving the overall quality of the graph data. Unlike existing methods, FedNGDB can handle complex graph structures and relationships, making it suitable for various downstream tasks. We evaluate FedNGDB on three real-world datasets, demonstrating its effectiveness in retrieving relevant information from multi-source graph data while keeping sensitive information secure on local devices. Our results show that FedNGDB can efficiently retrieve answers to cross-graph queries, making it a promising approach for large-scale LLMs and other applications that rely on efficient data retrieval mechanisms.
privacy preserving, neural graph databases, federated Learning
@@ 1. Introduction
Graph Databases (GDBs) are essential for efficiently storing and managing highly interconnected data in a graph structure. Their ability to handle complex relationship queries makes them invaluable for applications like recommendation systems (Wang et al., 2019; Cao et al., 2019) and fraud detection (Prusti et al., 2021; Sadowski and Rathle, 2014). GDBs offer flexibility with dynamic data models, while their performance and scalability ensure efficient query handling. In the era of large language models (LLMs), the significance of GDBs has grown, particularly with the Retrieval Augmented Generation (RAG) paradigm, where LLM agents utilize external GDBs like knowledge graphs (KGs) to enhance their retrieval capabilities (Gao et al., 2023; Lewis et al., 2020). This integration facilitates the creation of interactive natural language interfaces tailored to domain-specific applications, enabling more intuitive and accessible interaction with structured data and unlocking new possibilities for intelligent data-driven solutions (Matsumoto et al., 2024; Hussien et al., 2024; Jin et al., 2024). However, traditional graph databases often suffer from two limitations: the ineffectiveness of free text semantic search and graph incompleteness, which is a prevalent issue in real-world knowledge graphs and other graph-structured data. Incompleteness leads to the exclusion of relevant results, as the graph database may not capture all the necessary relationships and connections between entities by traversing (Bordes et al., 2013; Galárraga et al., 2013).
To address these limitations, neural graph databases (NGDBs) have recently been proposed (Ren et al., 2023; Besta et al., 2022). They integrate the adaptable structure of graph data models with the powerful processing capabilities of neural networks, allowing for the effective and efficient storage, and analysis of graph-structured data. NGDBs provide unified storage for diverse entries in an embedding space and neural query engine searching answers to input complex queries from the unified storage (Ren et al., 2023). These databases unlock stronger capabilities for intelligent data exploration, enabling users to craft complex queries and make informed inferences with the help of advanced neural network techniques. Among these applications, complex query answering (CQA) is an important yet challenging task in graph reasoning and can be used for supporting various downstream tasks (Bai et al., 2023c; Ren et al., 2023). CQA aims to retrieve answers that satisfy given logical expressions (Hamilton et al., 2018; Ren et al., 2020), which are often defined in predicate logic forms with relation projection operations, existential quantifiers $\exists$, logical conjunctions $\wedge$, disjunctions $\vee$, etc. As shown in Figure 1, given a logical query $q$, We aim to find all the research topic entities $V_{?}$ for which there exists Nobel Prize winners $V$ who was born in Germany and conducted studies in that specific field.
@@
While neural graph databases have achieved remarkable success in addressing complex query answering tasks, they are limited to utilizing a single central graph database and cannot be extended to multiple databases. As data assumes an increasingly vital role, NGDBs have experienced rapid growth in scale and scope, aggregating knowledge from diverse domains. Consequently, constructing a graph database that includes all related entities and relations has become difficult and it is impractical to access a central database with all the data needed (Peng et al., 2021a; Chen et al., 2021; Zhang et al., 2022). Collaborations between various NDGB holders are essential for answering more complicated queries. For example, as shown in Figure 1, there are multiple NGDBs with different domain knowledge. A complex query may consist of entities and relations from multiple NGDBs, preventing a local query answering model on a single database from answering that cross-graph queries. However, there are various reasons hindering the data sharing between NGDB holders. For example, the growing attention on privacy, regulations such as the General Data Protection Regulation (GDPR), and commercial interest between data holders, etc. To solve the challenge, some distributed databases have been proposed (Nadal et al., 2021), however, the study for the latest neural graph databases is still ignored.
@@
Figure 1. An example of cross graph queries on distributed neural graph databases. The relations and entities in a KG complex query can be from NGDBs which cannot be solved in a single local database.
@@
To solve the above challenges, federated learning (McMahan et al., 2017; Yang et al., 2019) has been proposed which allows multiple participants to distributedly train a global model collaboratively without raw data transmission. Federated learning has been widely applied in various domains (Long et al., 2020), such as knowledge graph embedding and federated databases. In such systems, raw graph triplets are kept on the local devices, participants train local models, and only gradient will be transferred to learn a global graph embedding model (Chen et al., 2021; Zhang et al., 2022; Peng et al., 2021a). In the training process, homomorphic encryption (HE) (Paillier, 1999), secure multi-party computation (SMPC) (Mohassel and Zhang, 2017), and differential privacy (Dwork, 2008; Geyer et al., 2017) are widely applied in federated learning to improve security. FedAvg (McMahan et al., 2017) is a commonly used technique in federated learning that updates global models by averaging local models trained on individual data and has been widely studied in various areas, including federated knowledge graph embedding (Chen et al., 2021; Zhang et al., 2022). Although federated learning can protect raw data, recent studies have indicated that learned representations can still potentially leak privacy (Hu and Song, 2023; Duddu et al., 2020), where an attacker can infer sensitive information from the learned embeddings.
@@
While federated learning has been widely applied in learning graph embeddings, existing works only focus on learning high-quality representations for simple downstream tasks, such as knowledge graph completion (Chen et al., 2021; Zhang et al., 2022; Huang et al., 2022; Tang et al., 2023), and lacks the ability to reason over graphs and retrieving answers to complex queries. To solve the constraints, we propose a Federated Neural Graph DataBase (FedNGDB), to reason over multi-source graphs avoiding sensitive raw data transmission to protect data privacy. FedNGDB can be applied to different central complex query answering models. It leverages federated learning techniques to train local query answering models in local NGDBs and align graph embeddings for global queries. Different from other federated graph embedding models, the FedNGDB server not only takes the responsibility of aggregating global models but also decomposing given complex queries to sub-queries to compute the query encoding and retrieve answers from distributed NGDBs under the protection of multi-party computation to avoid global model storage. Meanwhile, to better evaluate distributed NGDB systems’ performance, we create a benchmark on three widely used datasets. We evaluate our proposed FedNGDB on the benchmark and the experiment results show the effectiveness of retrieving answers to complex queries from multi-source graphs. We summarize our major contributions as follows:
@@
  • To the best of our knowledge, we are the first to extend federated graph embedding systems to complex query answering tasks, which is critical for graph holders’ collaboration.
@@
  • Based on three public datasets, we propose a benchmark for evaluating the retrieval performance of distributed NGDB systems. The benchmark systematically evaluates the retrieval performance facing cross-graph queries.
@@
  • We propose FedNGDB, a federated neural graph database system that can retrieve answers to complex queries from distributed NGDBs with privacy preserved. Extensive experiments conducted on three datasets demonstrate its high performance when facing cross-graph queries.
@@ 2. Related Work
2.1. Neural Graph Database
@@
Neural Graph Databases (NGDBs) neutralize traditional GDBs’ storage and query planning modules, aiming for stronger intelligent data exploration capabilities. Complex query answering (CQA) is a crucial task in NGDBs, which involves training a model to process and answer complex logical queries based on graph reasoning, a process known as query encoding. These methods represent complex queries into various structures and effectively search answers among candidate knowledge graph entities. GQE (Hamilton et al., 2018), Q2B (Ren et al., 2020) and HypeE (Liu et al., 2021) encode queries to vectors, hyper-rectangle and hyperbolic embeddings, respectively. To support negation operators, various encoding methods are proposed: Q2P (Bai et al., 2022) and ConE (Zhang et al., 2021) use multiple vectors to represent queries. BetaE (Ren and Leskovec, 2020), GammaE (Yang et al., 2022), PERM (Choudhary et al., 2021) propose to use various probabilistic distributions to encode complex logic graph queries. Some methods, like LMPNN (Wang et al., 2022) CQD (Arakelyan et al., 2020), Var2Vec (Wang et al., 2023a) take pre-train embeddings on simple link prediction tasks and apply logic operators to answer complex queries. Meanwhile, Neural structures are utilized to encode complex queries: BiQE (Kotnis et al., 2021) and KgTransformer (Liu et al., 2022) are proposed to use transformers, SQE (Bai et al., 2023c) applies sequential encoders, GNN-QE (Zhu et al., 2022) and StarQE (Alivanistos et al., 2021) use message-passing graph neural networks to encode queries, respectively. There are also encoding methods proposed to encode various knowledge graph types: NRN (Bai et al., 2023b) is proposed to encode numerical values, MEQE (Bai et al., 2023a) extends logical queries over events, states, and activities.
@@
While there are numerous existing complex query answering methods, these methods mainly focus on a single large graph. There are some methods proposed to reason over multi-view and temporal and varying graphs: MORA (Xi et al., 2022) ensembles multi-view knowledge graphs to scale up complex query answering, TTransE (Leblay and Chekol, 2018) and TRESCAL (Wang et al., 2015) can be applied on temporal knowledge graphs. However, these methods need raw data transmission and privacy protection is not considered. Conducting privacy-preserving complex query answering on multi-source knowledge graphs is still unexplored. With the growing attention on privacy and data protection, sensitive data cannot circulate freely among data holders, and complex query answering is forced to be conducted collaboratively on multiple knowledge graphs. Our research introduces federated learning to existing complex query answering so that we can apply reasoning over distributed NGDBs without raw data sharing.
2.2. Federated Databases
In recent years, federated learning has emerged as a promising approach to address privacy and scalability concerns in machine learning. It allows data owners to participate in model co-construction without raw data transmission to reduce privacy leakage risks (McMahan et al., 2017; Yang et al., 2019) under the protection of privacy protection techniques such as differential privacy (DP) (Yang et al., 2019), homomorphic encryption (HE) (Zhang et al., 2020), secure multi-party computation (SMPC) (Byrd and Polychroniadou, 2020). Various studies have been conducted to explore the potential of federated learning in different domains, such as recommender system (Yang et al., 2020), finance (Long et al., 2020) etc. Federated databases are proposed to manage distributed data management allowing storing and querying databases with privacy preserved (Bater et al., 2016; Berger and Schrefl, 2008; Li, 2013). Although Some federated graph databases are proposed to manage graph-based data (Nadal et al., 2021), the study for the latest neural graph databases are still be ignored.
@@ 2.2. Federated Databases
Federated knowledge graph embedding is another related topic. It tries to represent entities and their semantic relations into embedding spaces. FedE (Chen et al., 2021) learns knowledge graph embeddings locally and aggregates all local models in a global server for higher representation quality. FedR (Zhang et al., 2022) proposes to learn representation with privacy-preserving relation aggregation to avoid privacy leakage risks in entity embedding and reduce communication costs. FedCKE (Huang et al., 2022), FedMKGC (Tang et al., 2023) extend federated learning to learn global representations from different domains and multilingual knowledge graphs. FedEC (Chen et al., 2022b) applies contrastive learning to tackle data heterogeneity in knowledge graphs. MaKEr (Chen et al., 2022a) and MorseE (Chen et al., 2022c) utilize meta-learning to transfer knowledge among knowledge graphs to train graph neural networks for unseen knowledge extrapolation and inductive learning. DP-FLames (Hu et al., 2023) quantifies the privacy threats and incorporates private selection in federated knowledge embeddings. FLEST (Wang et al., 2023b) decomposes the embedding matrix and enables the sharing of latent representations to reduce the risks of privacy leakage and communication costs, FedM (Hu et al., 2022b) splits the duty of aggregating entities and relations to reduce the risks of graph reconstruction attacks. FKGE (Peng et al., 2021b) applies differential privacy and avoids the need for a central server. While existing federated knowledge graph embedding methods are proposed to distributedly learn high-quality representations with privacy preservation, there are some works indicate that the learned embeddings are informative and vulnerable to various privacy attacks (Hu et al., 2022a), even in federated scenarios (Hu and Song, 2023; Hu et al., 2023). Besides, these methods all lack the ability to answer complex queries on multi-source knowledge graphs which is critical for more complicated downstream tasks. Our research expands the simple federated knowledge graph embedding to answer complex queries on distributed knowledge graphs.